OpenStudy (anonymous):

the probability of winning a game is 15%. if you play 20 times, how many times should you expect to win? A. 5 times B. 3 times C. 6 times D. 15 times

OpenStudy (anonymous):

@xGuardians

geerky42 (geerky42):

I believe you just multiply 15% by 20

OpenStudy (anonymous):

that equal 300

geerky42 (geerky42):

That's because you do 15*20. You need to convert 15% to decimal.

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

so .15*20 = 3

OpenStudy (anonymous):

0.15

OpenStudy (pinksapphire):

Move the decimal point of the percent 2 places to the left so it will be 0.15

OpenStudy (anonymous):

ok you get 3

OpenStudy (pinksapphire):

Yes...

OpenStudy (pinksapphire):

Or you can do it this way...

OpenStudy (pinksapphire):

\[\frac{ 15 }{ 100 }=\frac{ x }{ 20 }\]

OpenStudy (pinksapphire):

Solve for x and it will equal 3.